OpenStudy (anonymous):

Two planets of rdii in the ratio of 2:3 are made from the materials of density in the ratio 3:2. Then the ratio of their accelerations due to gravity g1/g2 at the surface of the two planets will be

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (anonymous):

r is inversly propertional to d so r2/r1 = d1/d2

OpenStudy (stamp):

\[\frac{r_{1}}{r_{2}}=\frac{2}{3}\]\[\frac{d_{1}}{d_{2}}=\frac{3}{2}\]\[Find\ \frac{g_{1}}{g_{2}}\]

OpenStudy (anonymous):

g1/g2=3/2*2/3 =1

OpenStudy (anonymous):

i m not sure

OpenStudy (stamp):

\[g=\frac{Gm}{r^2},\ G=6.67^110^{-11}\]

OpenStudy (stamp):

\[d=\frac{m}{V}\]\[m=d*V\]

OpenStudy (anonymous):

u have confuse me

OpenStudy (stamp):

acceleration due to gravity is determined by the g(m,r) equation. G is some gravitational constant. What class is this for? Physics?

OpenStudy (shubhamsrg):

g = GM/r^2 M = VD V= 4/3 pi r^3 SO, does it help ?

OpenStudy (anonymous):

this is the question of class X

OpenStudy (anonymous):

but here are r1 and r2

OpenStudy (stamp):

r1 is the radius of the first planet, r2 is the radius of the second planet. d1 is the density of the first planet, d2 is the density of the second planet.